OpenStudy (anonymous):

Are 2x + 4 over 3x + 6 and 2over3 equivalent expressions? Using complete sentences explain why or why not.

OpenStudy (anonymous):

last question that i need help with, sorry thanks so much!

OpenStudy (diyadiya):

2x+4 = 2(x+2) 3x+6 = 3(x+2) 2(x+2)/3(x+2)= 2/3

OpenStudy (anonymous):

Thank you!